OpenStudy (anonymous):

On a set of house plans an architect represents a veranda of dimensions 7.2 m by 4.8 m by a rectangle 6 cm by 4 cm. Find the scale ratio used for the plans.

OpenStudy (anonymous):

Scale ratio is the actual size divided by the scale size. \[\frac{ actual size }{ scale size }= \frac{ 4.8 }{ 4 }=\frac{ 1.2 }{ 1 }\] So the scale ratio is 1.2 meters to 1centimeter. So 1 centimeter on the drawing equals 1.2 meters on veranda.